Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

probability dose for the day

Expert replies
by thephoenix » Mon Feb 22, 2010 10:51 pm
Mr. Yutaro’s class contains 5boys and 11 girls. If two students are chosen, one at a time, from the class, what is
the probability that a boy and a girl are chosen?
A. 9/48
B. 9/20
C. 11/24
D. 27/48
E. 13/16
Join the discussion
Source: — Problem Solving |

by ajith » Mon Feb 22, 2010 11:02 pm
thephoenix wrote:Mr. Yutaro’s class contains 5boys and 11 girls. If two students are chosen, one at a time, from the class, what is
the probability that a boy and a girl are chosen?
A. 9/48
B. 9/20
C. 11/24
D. 27/48
E. 13/16
No of ways in which 2 students can be chosen from the group of 16 (11+5) = 16C2 = 16*15/2 = 8*15 =120
No of ways in which 1 girl and 1 boy can be chosen from 11 girls and 5 boys = 11C1*5C1 = 55
[spoiler]
Probability = 55/120 = 11/24[/spoiler]
Always borrow money from a pessimist, he doesn't expect to be paid back.
Join the discussion

by analyst218 » Fri Feb 26, 2010 1:33 pm
Can neone help me understand this more clearly.
I understand why ajith is correct.
However, can't you also say

P(A) = boys
P(B) = girls

P(AandB) = choosing boys and girls?


P(A)*P(BgivenA) = [5/16]*[11/15] = 11/48

I get 11/24 as above when I do P(B)*P(B) = 11/16 * 10/15 = 11/24

Correct me what I'm thinking wrong.....

thanks
Join the discussion

by arzanr » Fri Feb 26, 2010 8:39 pm
analyst218 wrote:Can neone help me understand this more clearly.
I understand why ajith is correct.
However, can't you also say

P(A) = boys
P(B) = girls

P(AandB) = choosing boys and girls?


P(A)*P(BgivenA) = [5/16]*[11/15] = 11/48

I get 11/24 as above when I do P(B)*P(B) = 11/16 * 10/15 = 11/24

Correct me what I'm thinking wrong.....

thanks
Analyst,

You're thinking is correct but you stopped halfway. The question says that a boy and girl is picked which means that you can either first pick a boy and then a girl or first pick a girl and then a boy. You calculated the probability of picking a boy first and then a girl which is equal to 11/24. Now add the prob of picking the girl first and you get:

11/48 + 11/48 = 22/48 = 11/24
Join the discussion

by analyst218 » Mon Mar 01, 2010 11:17 am
picking boy and girl= p(boy then girl) + p(girl then boy)
= P(boy)P(girlgivenboy) + p(girl)p(boygivengirl)
= 11/48 + 11/48 = 11/24

makes total sense...
Join the discussion

by Brent@GMATPrepNow » Thu Jan 09, 2020 6:16 am
thephoenix wrote:Mr. Yutaro’s class contains 5boys and 11 girls. If two students are chosen, one at a time, from the class, what is
the probability that a boy and a girl are chosen?
A. 9/48
B. 9/20
C. 11/24
D. 27/48
E. 13/16
P(1 boy and 1 girl chosen) = P(1st selection is boy AND 2nd selection is girl OR 1st selection is girl AND 2nd selection is boy)
= P(1st selection is boy AND 2nd selection is girl) + P(1st selection is girl AND 2nd selection is boy)
= P(1st selection is boy) x P(2nd selection is girl) + P(1st selection is girl) x P(2nd selection is boy)
= 5/16 x 11/15 + 11/16 x 5/15
= 55/240 + 55/240
= 110/240
= 11/24
= C

Cheers,
Brent
Brent Hanneson - Creator of GMATPrepNow.com
Image
Join the discussion